### v1

@@
[datatype=bibtex]
[…]
A Systematic Approach to Predict the Impact of Cybersecurity Vulnerabilities Using LLMs
@@ Abstract
Index Terms:
vulnerability impact, CVE, Att&ck techniques, large language models, automated mapping.I Introduction
@@ II-C Automating the Mapping from CVE to ATT&CK
[16] [16] also apply SRL, and in contrast to SMET, which exclusively uses the attack vectors extracted from CVE descriptions, they also consider the surrounding context (the full CVE description) as input to the classification model. Moreover, they incorporate attack functionality prediction as an intermediate step towards predicting the attack techniques. This step partially draws on the CMM, as predicted functionalities can be mapped to ATT&CK techniques using the methodology’s predefined associations (method II-A-2). However, functionality mapping represents only one of the five methods outlined in the CMM, and it does not provide comprehensive coverage of all techniques. In contrast, our proposed Triage approach leverages the full range of methods from the CMM, along with labeled examples. Moreover, it is the only automated approach that explicitly models the three mapping types from the CMM (exploitation techniques, primary and secondary impacts), providing additional context to each mapping, and offering more complete insight on how a vulnerability is exploited.
@@
Finally, [21] [21] seem to explore an idea similar to ours, describing how they prompt ChatGPT based on the CMM and using few-shot examples. However, their exact contributions are difficult to evaluate, as their prompts only define overall mapping goals as the context and include a CVE description in the user prompt, omitting details for the five methods of the CMM, as well as the few-shot examples of labeled CVEs. The authors mention consolidating the primary and secondary impacts due to the difficulty of establishing the relationship and ordering between them.
@@ III The Hybrid Triage Approach
Our goal is to take a CVE as input and automatically map it to relevant attack techniques defined in Att&ck. We propose a hybrid approach with multiple separate LLM invocations to capture the various aspects of each CVE. Finally, we combine the output of individual runs to obtain a list of technique predictions per mapping type. Triage is illustrated in Figure 2 and consists of two main components, the Methodology Mappers, which implement rule-based mapping following the CMM, and the In-Context Learner, which implements many-shot in-context learning based on already labeled CVEs.
@@ III-A Methodology Mappers
As input, the model takes a CVE with its corresponding description. Using a template-based approach, we create prompts for each of the five methods of the CMM, and run these separately through the LLM. The prompt templates used for each method are detailed in Appendices -A to -E
@@ III-B In-Context Learner
The In-Context Learner creates attack mappings based on relevant features and in-context examples without any direct reference to the CMM. The prompt template is detailed in Section -F. It begins with categorizing an attack in three steps: Exploitation technique, Primary Impact and Secondary Impact. Next, we provide all enterprise attack techniques as classes to choose from. For each technique, we include the attack names and the full descriptions from Att&ck. Then, for each target CVE, we include CVSS scores and CWE types in addition to the CVE description. Next, we include in-context examples from our training set. Each in-context example includes CVE ID, CVE description, attack technique ID, attack name, and mapping type. The last part of the prompt summarizes the task and the output format, in which the LLM is instructed to predict a ranked list of the ten most relevant attack techniques for a given mapping type.
@@ V-C Ablation Study and Analysis of Classwise Predictions
(A in features are the attack technique descriptions;
#demo are the number of demonstrations included).
@@ V-F Addressing our Research Questions
There is a clear improvement in performance from our approach compared to SMET (Table V).
@@ VI Concluding Remarks
To support open science and allow for replication and verification of our work, a replication package with prompts, scripts, and results is made available. While under review, we share it through Google Drive [6] to allow for reviewer feedback. It will be deposited on Zenodo after acceptance.

### v2

@@ Abstract
I Introduction
@@ II-C Automating the Mapping from CVE to ATT&CK
[16] [16] also apply SRL to this task. However, in contrast to SMET, which exclusively uses the attack vectors extracted from CVE descriptions, they also consider the surrounding context (the full CVE description) as input to the classification model. Moreover, they incorporate attack functionality prediction as an intermediate step towards predicting the attack techniques. This step partially draws on the CMM, as predicted functionalities can be mapped to ATT&CK techniques using the methodology’s predefined associations (method II-A-2). However, functionality mapping represents only one of the five methods outlined in the CMM, and it does not provide comprehensive coverage of all techniques. In contrast, our proposed Triage approach leverages the full range of methods from the CMM, along with labeled examples. Moreover, it is the only automated approach that explicitly models the three mapping types from the CMM (exploitation techniques, primary and secondary impacts), providing additional context to each mapping, and offering more complete insight on how a vulnerability is exploited.
@@
Finally, [21] [21] explore an idea related to the present paper, describing how they prompt ChatGPT based on the CMM and few-shot examples. However, their contributions are difficult to evaluate, as their prompts only define overall mapping goals as the context and include a CVE description in the user prompt, omitting details for the five methods of the CMM, as well as the few-shot examples of labeled CVEs. The authors mention consolidating the primary and secondary impacts due to the difficulty of establishing the relationship and ordering between them.
@@ III The Hybrid Triage Approach
Our goal is to take a CVE as input and automatically map it to relevant attack techniques defined in Att&ck. We propose a hybrid approach with multiple separate LLM invocations to capture the various aspects of each CVE. Finally, we combine the output of individual runs to obtain a list of technique predictions per mapping type. Triage is illustrated in Figure 2 and consists of two main components, the Methodology Mappers, which implement rule-based mapping and prompting following the CMM, and the In-Context Learner, which implements many-shot in-context learning based on already labeled CVEs.
@@ III-A Methodology Mappers
As input, the model takes a CVE with its corresponding description. Using a template-based approach, we create prompts for each of the five methods of the CMM, and run these separately through the LLM. The prompt templates used for each method are detailed in Appendices -A to -E.
@@ III-B In-Context Learner
The In-Context Learner creates attack mappings based on relevant features and in-context examples without any direct reference to the CMM. The prompt was developed iteratively, adding relevant features and refining these based on the response. The prompt template is detailed in Section -F. It begins with categorizing an attack in three steps: Exploitation technique, Primary Impact and Secondary Impact. Next, we provide all enterprise attack techniques as classes to choose from. For each technique, we include the attack names and the full descriptions from Att&ck. Then, for each target CVE, we include CVSS scores and CWE types in addition to the CVE description. Next, we include in-context examples from our training set. Each in-context example includes CVE ID, CVE description, attack technique ID, attack name, and mapping type. The last part of the prompt summarizes the task and the output format, in which the LLM is instructed to predict a ranked list of the ten most relevant attack techniques for a given mapping type.
@@ V-F Addressing our Research Questions
There is a clear improvement in predictive performance for Triage compared to SMET (Table V).
@@
Comparing again with SMET, run-times or costs are unfortunately not reported by its authors, and since we do not train its models, we cannot compute these costs ourselves. However, when we run inference with SMET on an AMD EPYC 7413 CPU with 2.85 GHz, the processing time averages around 0.6 seconds per CVE. So, purely looking at run times, SMET outperforms Triage, although as we have seen, it does so at a lower predictive performance.
@@ V-G Threats to Validity
V-H Ethical considerations
We acknowledge that Triage has a measurable dual-use risk, considering that adversaries could use the predictions to prioritize exploitation paths or craft more effective payloads. Such dual-use risks are common for security-related tooling, e.g., a malware detector can be misused to assess sufficient obfuscation of malicious code. In the case of Triage, the risk is offset by substantial defensive benefits: by predicting the most likely Att&ck techniques for a vulnerability, it helps defenders prioritize patches, focus threat-hunting efforts, and map mitigations to realistic attack paths. As such, we believe that the benefits of providing defenders with capable tooling far outweigh the risks of misuse by attackers.
@@ VI Concluding Remarks
To support open science and allow for replication and verification of our work, a replication package with prompts, scripts, and results is made available via Zenodo [6].